OpenStudy (anonymous):

A family refers to two or more people related by marriage,birth, or adoption who reside together. in 2000, in Country A, the average networth of a family was $240,000 and there were about 8.8 * 10 to the seventh power families. Calculate the total networth of a family in 2000

OpenStudy (anonymous):

I came up with .0027 but I do not think that is right

jimthompson5910 (jim_thompson5910):

it says "in 2000, in Country A, the average networth of a family was $240,000 " so the question is asking what's already given...something seems off

OpenStudy (anonymous):

I typed it as it was written, I have to use scientific notation and round to the nearest tenth as needed \[\frac{ 2.4*10^{5} }{ ?8.8*10^{7} }\]

jimthompson5910 (jim_thompson5910):

i see how you got 0.0027, but the problem doesn't make sense. It gives you the net worth, but asks you what the net worth is.

OpenStudy (anonymous):

Then I came up with \[2.7\times10^{-2}\]

OpenStudy (anonymous):

I see what you are saying about that as well Im so lost lol

jimthompson5910 (jim_thompson5910):

\[\frac{ 2.4*10^{5} }{ 8.8*10^{7} } = \frac{2.4}{8.8} \times 10^{5-7}\] \[\frac{ 2.4*10^{5} }{ 8.8*10^{7} } = 0.27 \times 10^{-2}\] \[\frac{ 2.4*10^{5} }{ 8.8*10^{7} } = (2.7 \times 10^{-1}) \times 10^{-2}\] \[\frac{ 2.4*10^{5} }{ 8.8*10^{7} } = 2.7 \times 10^{-1-2}\] \[\frac{ 2.4*10^{5} }{ 8.8*10^{7} } = 2.7 \times 10^{-3}\]

jimthompson5910 (jim_thompson5910):

but I doubt that's what you're looking for since they want the net worth

OpenStudy (anonymous):

I am not sure where you came up with the 2.7 x \[10^{-1-2}\] because my examples from other problems just subtract straight down

jimthompson5910 (jim_thompson5910):

I'm adding the exponents because I'm multiplying 10^(-1) and 10^(-2)

jimthompson5910 (jim_thompson5910):

-1 + (-2) = -1 - 2

OpenStudy (anonymous):

No I get that part, what I do not get is where the \[10^{-1}\] even came from????

jimthompson5910 (jim_thompson5910):

oh 0.27 = 2.7 x 10^(-1)

OpenStudy (anonymous):

Well almost I am pretty sure that if you take the original equation and then

OpenStudy (anonymous):

divide the parts seperately it equals \[2.7\times10^{-2}\]

jimthompson5910 (jim_thompson5910):

the answer is 0.0027, which in scientific notation is 2.7 x 10^(-3)

jimthompson5910 (jim_thompson5910):

so you're missing that 10^(-1)

jimthompson5910 (jim_thompson5910):

well i mean \[\frac{ 2.4*10^{5} }{ 8.8*10^{7} } = 0.0027\]

OpenStudy (anonymous):

So Country A's total family net worth is .0027 dollars?

jimthompson5910 (jim_thompson5910):

problem is...that makes no sense whatsoever

jimthompson5910 (jim_thompson5910):

if that was true, then this is the poorest country ever (and I mean really poor)

OpenStudy (anonymous):

lol now you see my confusion

jimthompson5910 (jim_thompson5910):

yeah this problem is just messed up

jimthompson5910 (jim_thompson5910):

I'm sry

OpenStudy (anonymous):

maybe it is asking for the total networth of all the families combined.?

jimthompson5910 (jim_thompson5910):

hmm good point, but it says Calculate the total networth of A FAMILY in 2000 so idk

OpenStudy (anonymous):

I see the confusion. The problem is worded funny. It is your choice, finding the total seems more logical to me. As we are told in engineering, if you state your assumption you can not be wrong, haha. So state I am assuming they are asking for the total networth of all families. haha

OpenStudy (anonymous):

I don't like this problem ultimately.

jimthompson5910 (jim_thompson5910):

same here

OpenStudy (anonymous):

Is it maybe asking that i use the scientific notation of simplifying the $240,000 down to 2.4 x \[10^{5}\]

jimthompson5910 (jim_thompson5910):

I'd just say the answer is $240,000...that's what they're asking for

jimthompson5910 (jim_thompson5910):

sure you can write the answer in sci notation if you want

OpenStudy (anonymous):

Thank you to you both, I'm simply going to show the work and then explain where the word problem became confusing as to what it was asking for.

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

The problem does not make meaning something is missing

OpenStudy (anonymous):

That is the entire problem. and then it says Country A's total family net worth is ____ dollars? (use scientific notation, and round to the nearest tenth as needed

OpenStudy (anonymous):

well the answer ended up being \[2.1\times10^{13}\]

jimthompson5910 (jim_thompson5910):

hmm there's probably some formula used to get that...not sure which one